i need help with this one problem. Find three consecutive even integers such that three times the middle integer is six more than the summ of the first and third.
Answer by ankor@dixie-net.com(12704)

Just write an expression or equation for what it says:
"Find three consecutive even integers"
x, (x+2), (x+4)
:
" such that three times the middle integer is six more than the sum of the first and third."
3(x+2) = 6 + x + (x+4)
:
3x + 6 = 2x + 10
:
3x - 2x = 10 - 6
:
x = 4 is the 1st integer, then 6, 8 are the next two
:
:
Check solution in the statement:
"three times the middle integer is six more than the sum of the first and third."
3(6) = 6 + 4 + 8; confirms our solution
:
Not that hard, right?